OpenStudy (anonymous):

could someone please simply 5/14.5??? i would help them out with a problem :)

OpenStudy (anonymous):

i mean simplify sorry

OpenStudy (anonymous):

Divide by 5 on the top and bottom: 1/2.9 Or you could switch that to improper fractions and play with it a little: 14.5= 14+ (1/2) = 29/2, so 5/14.5 = 5/(29/2) Multiply by 2 on the top and the bottom = 10/29 You can leave it like this, because it's not normal to have decimal notations in fractions, or you could divide by 10 on top and bottom: = 1/2.9